OpenStudy (lulubj):

which is the standard equation for a circle centered at the origin with radius r? A. x^2+y^2=r^2 B. x^2=y^2+r^2 C. x^2+y^2=r D. x+y=r

OpenStudy (solomonzelman):

The (vertex form of the) equation of a circle is: \(\large\color{black}{ \displaystyle (x-{\rm \color{red}{h}} )^2+ (y-{\rm \color{blue}{k}} )^2 = {\rm \color{green}{r}}^2 }\) Where 1) \(\large\color{black}{ \displaystyle({\rm \color{red}{h}} ,~{\rm \color{blue}{k}} )}\) is the \(\normalsize\color{red}{ \bf c\color{blue}{e}n\color{blue}{t}e\color{blue}{r} }\) of the circle. 2) \(\large\color{black}{ \displaystyle {\rm \color{green}{r}} }\) is the \(\normalsize\color{green}{ \bf radius }\) of the circle.

OpenStudy (lulubj):

So A?

OpenStudy (solomonzelman):

yup.

OpenStudy (lulubj):

Thanks

OpenStudy (solomonzelman):

I was about to give you a tip to get to that answer, but no need in that.... good job!